The mean grade in a Physics examination was 75 and the standard deviation was 5. Determine the standard scores (z-value) corresponding to a grade of 85


1
Expert's answer
2021-11-11T12:29:19-0500

"\mu=75 \\
\sigma = 5 \\
x = 85 \\
Z = \frac{x - \mu}{\sigma} = \frac{85-75}{5} = 2"